Grade 11Mechanics

A ball is dropped from a height of 16 m on a hard ground and after collision it bounces back to a height h metre. If coeffiecient of restitution=3/4 and g=10m/s2. then find value of h

Aditya Saini
9 Years agoGrade 11

1 Answer

DR STRANGE
ApprovedApproved Tutor Answer9 Years ago
Hi Aditya ,
r2= h2/h1 , 9/16=h2/16 ,h2=9m
if you want derivation of this formula you can ask :)